# Learnable Koopman-Enhanced Transformer-Based Time Series Forecasting with Spectral Control

### v1

@@ 1 Introduction
Time series forecasting is a foundational task across science and engineering, underpinning applications in retail demand prediction [1, 2, 3], traffic flow management [4, 5], energy balancing [6, 7, 8], and financial volatility modeling [9, 10]. The increasing scale, dimensionality, and non-stationarity of modern datasets challenge traditional statistical models [11], driving the development of deep learning architectures that offer improved expressiveness, scalability, and feature abstraction.
@@
To improve scalability and accuracy, numerous specialized Transformers have been proposed. Informer [21, 16] introduces ProbSparse attention to focus on the most informative queries; Autoformer [22, 15] integrates decomposition blocks to explicitly handle trend and seasonal patterns; FEDformer [23, 24] employs Fourier spectral mixing to reduce complexity; and Pyraformer [25, 26] introduces a pyramidal hierarchy to capture multi-scale structures. More recently, PatchTST [17] showed that simple architectural principles—patching, channel-independence, and lightweight attention—are sufficient to outperform many deeper Transformer variants [27, 28].
Following the surprising results of DLinear [28, 29], which showed that a channel-wise linear model can outperform many Transformers on standard benchmarks, a central question has emerged: how much architectural complexity is actually necessary for strong time-series forecasting performance? Moreover, deep models still suffer from limited interpretability, sensitivity to distribution shifts [30, 31, 32], and a lack of explicit dynamical structure, limiting their robustness in high-stakes settings.
@@ 1.1 Koopman Operators and Deep Learning
The Koopman operator offers a linear lens for analyzing nonlinear dynamical systems by evolving observables rather than states. Formally, it is an infinite-dimensional linear operator acting on functions of the state, enabling the use of linear spectral tools in settings that may be nonlinear, nonstationary, or even chaotic [37, 38]. Classical data-driven approaches such as Dynamic Mode Decomposition (DMD) approximate the Koopman spectrum for prediction and modal analysis [39, 40], but tend to degrade under high-dimensional, noisy, or multivariate real-world data. Dynamic Mode Decomposition (DMD) offer practical means to approximate the operator from measurements and facilitate modal analysis, prediction, and control in complex systems [41, 42, 43].
@@
Compared to existing Koopman-based forecasting models [47, 50, 48], DeepKoopFormer is the first to offer provable spectral decay, Lyapunov-stable propagation, normality of the transition matrix, and closed-form error control—all within a general, extensible encoder–propagator–decoder pipeline aligned with the goals of physics-informed forecasting [53, 54] and interpretable dynamical systems modeling [55, 56].
@@ 1.2 Contributions: Learnable Koopman–Transformer Framework
We conduct an extensive simulation benchmark comparing the proposed learnable Koopman variants against two reference Koopman baselines—a spectrally constrained variant and an unconstrained variant—together with widely used forecasting models including LSTM, DLinear [28, 29], and a diagonal State-Space Model (SSM) [57]. Alongside forecasting accuracy, we provide the first systematic spectral evaluation of Koopman–Transformers, including density characterization, eigenvalue training trajectories, stability envelopes, and horizon sensitivity.
@@
Compared to existing Koopman-based forecasting models [47, 50, 48], Learnable-DeepKoopFormer is the first to offer provable spectral decay, Lyapunov-stable propagation, normality of the transition matrix, and closed-form error control—all within a general, extensible encoder–propagator–decoder pipeline aligned with the goals of physics-informed forecasting [53, 54] and interpretable dynamical systems modeling [55, 56].
@@ 3.1.6 Invertibility in finite dimensions and Banach space extension
In contrast, standard state-space models with unconstrained transition matrices $A$ (as commonly used in classical linear SSMs and recent deep state-space architectures) provide no such guarantees: $A$ may be singular, ill-conditioned, or unstable, and even when invertible its spectral properties are not explicitly controlled [61, 62, 63].
@@ 4 Numerical Simulations
For climate analysis and renewable energy applications, we employ atmospheric data from CMIP6555https://cds.climate.copernicus.eu/datasets/projections-cordex-domains-single-levels?tab=overview and ERA5666https://cds.climate.copernicus.eu/datasets, capturing regional wind speed and surface pressure over Germany [64, 65, 8]. As a complementary nonlinear system, we analyze cryptocurrency market behavior using a publicly available multivariate dataset [^7] https://github.com/Chisomnwa/Cryptocurrency-Data-Analysis, which includes volatility, pricing trends, and trading activity for major digital assets—an archetypal example of stochastic, regime-switching dynamics. Finally, we examine trends in national energy production by modeling electricity generation in Spain [^8] https://github.com/afshinfaramarzi/Energy-Demand-electricity-price-Forecasting/tree/main across fossil, wind, solar, hydro, and biomass sources, where physical constraints and weather dependencies induce complex temporal couplings. Together, these datasets span distinct dynamical regimes and provide a rigorous foundation for assessing the ability of Koopman-enhanced representations to extract latent linear structure from nonlinear temporal processes.

### v2

@@ 1 Introduction
Time series forecasting is a foundational task across science and engineering, underpinning applications in retail demand prediction [7, 41, 31], traffic flow management [26, 20], energy balancing [8, 32, forootani_cadnn_2025], and financial volatility modeling [5, 44]. The increasing scale, dimensionality, and non-stationarity of modern datasets challenge traditional statistical models [14], driving the development of deep learning architectures that offer improved expressiveness, scalability, and feature abstraction.
@@
To improve scalability and accuracy, numerous specialized Transformers have been proposed. Informer [47, 46] introduces ProbSparse attention to focus on the most informative queries; Autoformer [autoformer, 43] integrates decomposition blocks to explicitly handle trend and seasonal patterns; FEDformer [fedformer, 48] employs Fourier spectral mixing to reduce complexity; and Pyraformer [pyraformer, 22] introduces a pyramidal hierarchy to capture multi-scale structures. More recently, PatchTST [30] showed that simple architectural principles—patching, channel-independence, and lightweight attention—are sufficient to outperform many deeper Transformer variants [multichannel, dlinear].
Following the surprising results of DLinear [dlinear, zeng2023transformers], which showed that a channel-wise linear model can outperform many Transformers on standard benchmarks, a central question has emerged: how much architectural complexity is actually necessary for strong time-series forecasting performance? Moreover, deep models still suffer from limited interpretability, sensitivity to distribution shifts [18, 23, 16], and a lack of explicit dynamical structure, limiting their robustness in high-stakes settings.
@@ 1.1 Koopman Operators and Deep Learning
The Koopman operator offers a linear lens for analyzing nonlinear dynamical systems by evolving observables rather than states. Formally, it is an infinite-dimensional linear operator acting on functions of the state, enabling the use of linear spectral tools in settings that may be nonlinear, nonstationary, or even chaotic [Mezic2017book, khosravi2023representer]. Classical data-driven approaches such as Dynamic Mode Decomposition (DMD) approximate the Koopman spectrum for prediction and modal analysis [42, Kutz2016book], but tend to degrade under high-dimensional, noisy, or multivariate real-world data. Dynamic Mode Decomposition (DMD) offer practical means to approximate the operator from measurements and facilitate modal analysis, prediction, and control in complex systems [10, 34, 9].
@@
Compared to existing Koopman-based forecasting models [24, 37, 2], DeepKoopFormer is the first to offer provable spectral decay, Lyapunov-stable propagation, normality of the transition matrix, and closed-form error control—all within a general, extensible encoder–propagator–decoder pipeline aligned with the goals of physics-informed forecasting [15, 29] and interpretable dynamical systems modeling [Brunton2016plosone, Korda2016arxiv].
@@ 1.2 Contributions: Learnable Koopman–Transformer Framework
We conduct an extensive simulation benchmark comparing the proposed learnable Koopman variants against two reference Koopman baselines—a spectrally constrained variant and an unconstrained variant—together with widely used forecasting models including LSTM, DLinear [dlinear, zeng2023transformers], and a diagonal State-Space Model (SSM) [12]. Alongside forecasting accuracy, we provide the first systematic spectral evaluation of Koopman–Transformers, including density characterization, eigenvalue training trajectories, stability envelopes, and horizon sensitivity.
@@
Compared to existing Koopman-based forecasting models [24, 37, 2], Learnable-DeepKoopFormer is the first to offer provable spectral decay, Lyapunov-stable propagation, normality of the transition matrix, and closed-form error control—all within a general, extensible encoder–propagator–decoder pipeline aligned with the goals of physics-informed forecasting [15, 29] and interpretable dynamical systems modeling [Brunton2016plosone, Korda2016arxiv].
@@ 3.1.6 Invertibility in finite dimensions and Banach space extension
In contrast, standard state-space models with unconstrained transition matrices $A$ (as commonly used in classical linear SSMs and recent deep state-space architectures) provide no such guarantees: $A$ may be singular, ill-conditioned, or unstable, and even when invertible its spectral properties are not explicitly controlled [gu2021s4, smith2022simplified, hamilton2020time].
@@ 4 Numerical Simulations
For climate analysis and renewable energy applications, we employ atmospheric data from CMIP6555https://cds.climate.copernicus.eu/datasets/projections-cordex-domains-single-levels?tab=overview and ERA5666https://cds.climate.copernicus.eu/datasets, capturing regional wind speed and surface pressure over Germany [makula2022coupled, hersbach2020era5, forootani_cadnn_2025]. As a complementary nonlinear system, we analyze cryptocurrency market behavior using a publicly available multivariate dataset [^7] https://github.com/Chisomnwa/Cryptocurrency-Data-Analysis, which includes volatility, pricing trends, and trading activity for major digital assets—an archetypal example of stochastic, regime-switching dynamics. Finally, we examine trends in national energy production by modeling electricity generation in Spain [^8] https://github.com/afshinfaramarzi/Energy-Demand-electricity-price-Forecasting/tree/main across fossil, wind, solar, hydro, and biomass sources, where physical constraints and weather dependencies induce complex temporal couplings. Together, these datasets span distinct dynamical regimes and provide a rigorous foundation for assessing the ability of Koopman-enhanced representations to extract latent linear structure from nonlinear temporal processes.